The measure of the supplement of an angle exceeds three times the measure of the complement of the angle by 12. Find the measure of half of the supplement.
Answer by stanbon(75887)

The measure of the supplement of an angle exceeds three times the measure of the complement of the angle by 12. Find the measure of half of the supplement.
------------------------------------
Let the angle be "x".
Supplement is 180-x
complement is 90-x
---------------------------
Equation:
180-x - 3(90-x) = 12
180-x - 270 + 3x = 12
2x = 102
x = 51 degrees
---------------------
The supplement is 180-51 = 129 degrees
---
Final answer: half the supplement = 129/2 = 64.5 degrees
==============================================================
Cheers,
Stan H.